OpenStudy (anonymous):

Let a,b,c be +ve integers. If b=sqrt(ac), which of the following must be true? A. loga^2, logb^2, logc^2 is an arithmetic sequence. B. a^3, b^3, c^3 is a geometric sequence. C. 4^a, 4^b, 4^c is a geometric sequence.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

b= sqrt(ac) means a,b,c are a geometric sequence. remember yesterday ? \[ \frac{a}{b} = \frac{b}{c} \\ b^2= ac \\ b= \sqrt{ac} \]

OpenStudy (phi):

so replace a,b,c with a, k a, k^2 a and put those values into each of the choices and see what happens

OpenStudy (anonymous):

logb^2=log(ac) ....

OpenStudy (phi):

for choice A, replace a with a, b with k a, and c with k^2 a you get loga^2, logb^2, logc^2 --> log( a^2), log( (ka)^2), log( (k^2a)^2 ) remember the rules \[ \log(a^b) = b \log(a) \\ \log(xy) = \log(x)+ \log(y) \] can you simplify log( a^2), log( (ka)^2), log( (k^2a)^2 ) ?

OpenStudy (anonymous):

2loga 2logk+2loga 4a logk it seems quite strange...

OpenStudy (phi):

the third term becomes \[ \log\left( (k^2 a)^2 \right)= 2 \log(k^2 a)= 2(\log(k^2) + \log(a))=\\ 4 \log(k)+ 2\log(a) \]

OpenStudy (anonymous):

ohh...

OpenStudy (phi):

so you have the sequence 2loga 2logk+2loga 4logk +2loga find the difference between terms. what do you get ?

OpenStudy (anonymous):

2logk

OpenStudy (phi):

so you have a common difference of 2 logk so you have an Arith. seq.

OpenStudy (phi):

for B and C, look at the problems you did yesterday, because I think we already did these.

OpenStudy (anonymous):

B is correct

OpenStudy (phi):

or, you can just work it through. replace a,b,c with a, a k and a k^2 ...

OpenStudy (phi):

B is a G.P.

OpenStudy (anonymous):

a^3=a^3 b^3=k^3 a^3 c^3=k^6 a^3 common ratio is k^3, isn't it...

OpenStudy (anonymous):

C is also correct

OpenStudy (phi):

yes, B is true, with a common ratio of k^3 b^3/a^3 is (k^3 a^3)/a^3= k^3

OpenStudy (phi):

C is not true. How did you do it ?

OpenStudy (anonymous):

4^a=4^a 4^b=4^(ka) 4^c=4^(k^2 a) common ratio is .... i don't know how to express...

OpenStudy (anonymous):

why not?

OpenStudy (phi):

the ratios are 4^b / 4^a is 4^(b-a) using the rule \[ \frac{a^x}{a^y} = a^x \cdot a^{-y} = a^{x-y} \] and 4^c/4^b = 4^(c-b) if we replace a,b,c we will see that these ratios will be different

OpenStudy (anonymous):

oic

OpenStudy (anonymous):

thx

OpenStudy (phi):

we did a version of choice C yesterday